Solve :
-4.2 + 0.02x - 0.4 = 2y-5

Answer by Alan3354(69443)

Solve :
-4.2 + 0.02x - 0.4 = 2y-5
There are 2 unknowns, x and y, and only 1 equation. No single solution is possible.
Some of the terms can be collected, tho.
-4.2 + 0.02x - 0.4 = 2y-5
-4.6 + 0.02x = 2y-5
0.4 + 0.02x = 2y
The fractions can be eliminated:
20 + x = 100y
That's all.